Differentiate F(x)= arc sinx . sinx²


1
Expert's answer
2020-04-13T14:45:05-0400

Consider the function "F(x)=\arcsin(x)\cdotp \sin(x^2)"

Use product rule to differentiate the function as follows:


"F'(x)=\frac{d}{dx}[\arcsin(x)\cdotp \sin(x^2)]"


"F'(x)=\arcsin(x)\frac{d}{dx}(\sin(x^2))+(\sin(x^2))\frac{d}{dx}(\arcsin(x))"

Using chain rule, the derivative of "\sin(x^2)" with respect to "x" is evaluated as,


"\frac{d}{dx}(\sin(x^2))=\cos(x^2)\cdotp \tfrac{d}{dx}(x^2)=2x\cos(x^2)"

The derivative of "\arcsin(x)" with respect to "x" is evaluated as,


"\tfrac{d}{dx}(\arcsin(x))=\tfrac{1}{\sqrt{1-x^2}}"

Now, substitute the derivatives to obtain,


"F'(x)=\arcsin(x)\cdotp2x\cos(x^2)+(\sin(x^2))\cdotp\tfrac{1}{\sqrt{1-x^2}}"



Therefore, the derivative of the function is

"F'(x)=2x\cos(x^2)\cdotp\arcsin(x)+\tfrac{\sin(x^2)}{\sqrt{1-x^2}}"